OpenStudy (amonoconnor):

I need help figuring out this one... From square one, I just don't know what the heck it's talking about, but beyond that, how does one solve this analytically? The problem: "A rectangle has its base on the x-axis and its upper two vertices on the parabola 'y = 12 - x^2'. What is the largest area the rectangle can have?" Any and all help is greatly appreciated!! :)

OpenStudy (amonoconnor):

Hey @wolfe8, you up for a little Calculus? ;)

OpenStudy (wolfe8):

Can I give you a link that might help? http://answers.yahoo.com/question/index?qid=20090404100038AATDsJu

OpenStudy (amonoconnor):

. . . god, I do love yahoo answers, but I'm using a computer issued by my school which unfortunately has internet blocks to suspect blogging/ social sites. Sorry! But no... I can 'to use that :/

OpenStudy (amonoconnor):

Can you copy and paste text from the we page into a reply?

OpenStudy (wolfe8):

Sure. And sorry in advance for not going through it yourself. I would if my brain would cooperate. The parabola is symmetric around x=0. Therefore, if (x,12-x²) is one of the vertices of the rectangle, so is (-x, 12-x²). Therefore, the width of the rectangle is x - (-x) = 2x and its height is 12-x². You want to maximize the area A = 2x(12-x²) so take a derivative and set it equal to 0: 24-6x² = 0 or x²=4 leading to the solution of x=2. Thus, the width is 4, the height is 8, and the area is 32.

OpenStudy (wolfe8):

myself*

OpenStudy (amonoconnor):

It's okay :) I understood ;)

OpenStudy (amonoconnor):

How, analytically, do we know that the graph is symmetric around the Y-axis?

OpenStudy (wolfe8):

Perhaps by using the vertex equation?

OpenStudy (amonoconnor):

And what're the base equations used? What did they take the derivative OF? I think I get the concept they're shooting at here, but what's the work? Can I seee some algebra? That's something I can fully conceptualize and understand. I know it can be tedious to type out, but it would really help me out! :)

OpenStudy (ybarrap):

The height of the rectangle is can be found by taking x=0. The width of the rectangle can be found by taking y =0.  So Area = 12 * 2*sqrt(12)